OpenStudy (anonymous):

The shortest distance from the curve xy = 4 to the origin is

myininaya (myininaya):

i did this one early for someone else let me see if i can link you to it

OpenStudy (anonymous):

Thanks. anyone know this one: If f(x) = I(x^2 - 12)(x^2+4)I, how many numbers in the interval [-2,3] satisfy the conclusion of the mean value theorem?

myininaya (myininaya):

i cant find it lol anyways y=4/x \[d=\sqrt{(x-0)^2+(y-0)^2}=\sqrt{x^2+(\frac{4}{x})^2}\] \[d^2=x^2+\frac{16}{x^2}\]

myininaya (myininaya):

to minimize d all we need to do is minimize d^2

OpenStudy (anonymous):

I sort of get confused when minimizing

OpenStudy (anonymous):

hey myinin could you use legrange multipliers?

OpenStudy (zarkon):

what would your constraint be?

myininaya (myininaya):

\[(d^2)'=2x-2*16*\frac{1}{x^3}=\frac{2x*x^3}{x^3}-32\frac{1}{x^3}=\frac{2x^4-32}{x^3}\] \[=2*\frac{x^4-16}{x^3}\] critical numbers are -2,2,0 except f(0) doesn't exist so we just need to see which gives us shortest distance x=2 or x=-2 and they both give shortest distance

OpenStudy (anonymous):

x^2+y^2

OpenStudy (zarkon):

that is not a constraint

myininaya (myininaya):

(2,4/2) and (-2,-4/2)

myininaya (myininaya):

whats a constraint?

OpenStudy (zarkon):

xy=4 would be a constarint

myininaya (myininaya):

y=4/x

OpenStudy (zarkon):

there is a really simple way to do this problem

OpenStudy (anonymous):

ok sorry yeah lamda(del(constarint)=x^2+y^2

OpenStudy (anonymous):

del(x^2+y^2)

myininaya (myininaya):

wait zarkon i can't do that way?

OpenStudy (zarkon):

your way is fine

OpenStudy (zarkon):

I have a quick way to do it

myininaya (myininaya):

show and tell please

OpenStudy (anonymous):

I dont really know anything about constraints... myininaya, how did you get that equation for the derivative of d^2?

myininaya (myininaya):

i square both sides of the distance formula

myininaya (myininaya):

do you know how i got d right?

myininaya (myininaya):

hey zarkon you show me the quick way in a different thread i guess if trap doesn't want to see

OpenStudy (anonymous):

min google lagrange multipliers thats an easy way to do it

OpenStudy (anonymous):

I'd like to see!

myininaya (myininaya):

me too im so excited

OpenStudy (zarkon):

\[xy=4\] so \[y=\frac{4}{x}\] then \[y'=-\frac{4}{x^2}\] the shortest distance is the line that is perpendicular to the tangent this will happen when the the slope of the line from the origin to the function is 45deg...ie has slope 1 so solve \[-1=-\frac{4}{x^2}\] \[x^2=4\] staying in the first quadrant x=2

OpenStudy (zarkon):

plug the 2 in the distance formula

OpenStudy (zarkon):

this works really well when the function is really complicated. This function was fairly simple so it only helps a little.

OpenStudy (anonymous):

creative i like it zarkon

myininaya (myininaya):

this is a stupid question but where does -1 come from -4/x^2=tan(pi/4) ?

OpenStudy (anonymous):

where did the -1 come in?

OpenStudy (zarkon):

the slope of a perpendicular line is the negative reciprocal on the original line

OpenStudy (zarkon):

our slope was 1 so the perpendicular slope is -1/1=-1

myininaya (myininaya):

ok!

OpenStudy (zarkon):

can anyone do it with Lagrange multipliers :)

myininaya (myininaya):

nut was talking about lagrange multipliers i'm not sure

myininaya (myininaya):

zarkon how can i be smart like you?

OpenStudy (zarkon):

\[\nabla F=\lambda \nabla G\]

OpenStudy (zarkon):

you are smart

myininaya (myininaya):

not as smart

OpenStudy (zarkon):

I have a lot of experience. I've been doing this stuff for many years

myininaya (myininaya):

so the older i get hopefully the wiser i will get

OpenStudy (anonymous):

so 2 has to be plugged into the distance formula? It is not the answer?

OpenStudy (anonymous):

so 2 has to be plugged into the distance formula? It is not the answer?

OpenStudy (zarkon):

\[F(x,y)=x^2+y^2\] \[G(x,y)=xy-4\]

OpenStudy (zarkon):

2 is not the answer

OpenStudy (zarkon):

you have to plug it into the distance formula

myininaya (myininaya):

asking for shortest distance to find distance you need to plug 2 into distance formula since this will give you a distance

OpenStudy (anonymous):

what do I use for the y value?

OpenStudy (zarkon):

\[\frac{\partial F}{\partial x}=2x\] \[\frac{\partial F}{\partial y}=2y\] \[\frac{\partial G}{\partial x}=y\] \[\frac{\partial G}{\partial y}=x\]

myininaya (myininaya):

remember i gave you y earlier f(2)=4/2=2 so use (2,2) and (0,0) to find distance

OpenStudy (zarkon):

\[2x=\lambda y\] \[2y=\lambda x\]

OpenStudy (zarkon):

\[\frac{2x}{y}=\frac{2y}{x}\Rightarrow x^2=y^2\]

OpenStudy (zarkon):

restrict to 1st quadrant

OpenStudy (zarkon):

xy=4 xx=4 x=2

myininaya (myininaya):

lol so are there anymore ways?

OpenStudy (anonymous):

thanks! would anyone be able to hep me with the other problem i posted in this thread?

OpenStudy (zarkon):

good stuff...so many ways to do it

OpenStudy (zarkon):

\[f(x) = |(x^2 - 12)(x^2+4)|\] \[f(x)=(12-x^2)(x^2+4)\] on [-2,3]

OpenStudy (anonymous):

how did you switch around thhose numbers?

OpenStudy (zarkon):

on [-2,3] x^2-12 is negative

OpenStudy (zarkon):

so \[|x^2-12|=|-(12-x^2)|=|12-x^2|=12-x^2\]

OpenStudy (zarkon):

on [-2,3]

OpenStudy (zarkon):

\[x^2+4\] is always positive

OpenStudy (zarkon):

was it a product or division?

myininaya (myininaya):

oops

OpenStudy (zarkon):

I get 3 numbers

myininaya (myininaya):

lol

OpenStudy (zarkon):

I cheated to get my answer though ;)

myininaya (myininaya):

i blame that it is getting later

myininaya (myininaya):

late*

OpenStudy (zarkon):

you on the east coast?

myininaya (myininaya):

central its almost 1:00 am

OpenStudy (zarkon):

Arkansas?

myininaya (myininaya):

yes

OpenStudy (anonymous):

how did you cheat? I applied the MVT and I'm stuck with an equation I dont know how to solve.

OpenStudy (zarkon):

I wrote a program on my calculator a while back that tells me the naswer

OpenStudy (anonymous):

hey can someone check this log derivative for me, only check work please?

OpenStudy (anonymous):

hey can someone check this log derivative for me, only check work please?

myininaya (myininaya):

what equation did you get trapp?

OpenStudy (anonymous):

-x^4 + 8x^2 + 48 = -5. I played around with it, but...

OpenStudy (zarkon):

my calculator could not give me exact answers

OpenStudy (zarkon):

it gave me decimal approximations

OpenStudy (anonymous):

this section is noncalc.

OpenStudy (zarkon):

that is your derivative?

myininaya (myininaya):

i got f'=-4x^3+16x

OpenStudy (zarkon):

that's correct

OpenStudy (anonymous):

oooh I forgot to take the deriv after i simplified ok hold on

myininaya (myininaya):

this doesn't become a pretty equation

OpenStudy (zarkon):

no

OpenStudy (anonymous):

ok got that. so set equal to -5?

myininaya (myininaya):

no i set equal to -10/65

OpenStudy (anonymous):

how did you get that number?